# Drug Discovery SMILES-to-Pharmacokinetics Diffusion Models with Deep Molecular Understanding

### v1

@@ 3.5 Data Processing
We first merge all 10 PK datasets to create a unified dataset containing 30K drugs over 12 unique PK columns for training and testing (90%/10% split) our models. Excluding the hERG dataset from which we sample 7.9K drugs, we merge the remaining 9 PK datasets for 22.1K unique drugs. We arrive at a total of 30K drugs in our unified dataset after merging the 7.9K drugs sampled from hERG into our 22.1K unique drugs from the other 9 PK datasets. We only sample 7.9K drugs from hERG to maintain balance in the unified dataset given the size imbalance of hERG compared to the other 9 PK datasets. After removing outliers ($Q1-1.5$IQR lower and $Q3+1.5$IQR upper bound), we are left with 28,397 drugs from the original 30K drugs. The 28,397 drug values for each of the 12 PK columns are then min-max scaled between the range of $[-1,1]$. Before infilling null values using one of the average, uniform or Gaussian distributions, or the proposed DLGN method, we store the null masks for each drug for the masked loss function.
4 Experiments
Below we describe model training details and compare our synthetic data to real data, in terms of machine learning efficiency (MLE) and univariate and bivariate statistical distributions. We then discuss ablation studies and key findings. Metrics for MLE, univariate, and bivariate evaluations are further defined in their respective subsections.
4.1 Training Details
Imagand Model Diffusion Training
Layers 12 Learning Rate 1e-3
Heads 16 Weight Decay 5e-2
MLP Dim. 768 Epoch 3000
Emb. Dropout 10% Batch Size 256
Num Patches 48 Warmup 200
Cond. Emb. Size 768 Timesteps (Train) 2000
Time Emb. Size 64 Timesteps (Infer.) 150
PK Emb. Size 256 EMA Gamma ($\gamma$) 0.994
@@ 4.1 Training Details
4.2 Machine Learning Efficiency
Machine Learning Efficiency (MLE) is a measure that assesses the ability of the synthetic data to replicate a specific use case [10, 6, 7]. MLE represents the ability of the synthetic data to replace or augment real data in downstream use cases. To measure MLE, two models are trained separately using synthetic versus real data, and then their performance, measured by mean-squared error, R2, and Pearson correlation coefficient, is evaluated on real data test sets and compared.
Figure 3: Synthetic PK Data Hellinger Distances (HDs).
Table 1 shows the results of the PK regression tasks using real and synthetic augmented datasets. Results of these experiments suggest that a synthetic augmented dataset can outperform real data with statistical significance over many PK datasets. Additional tasks will be explored in future work.
4.3 Univariate Distributions
@@
Table 3 compares the mean and standard deviation of the real and synthetic target property values. The mean and standard deviation of the generated synthetic data closely resemble that of the real data for each PK target property. We found that normalization combined with static thresholding substantially limits the generation of invalid and out-of-range PK values.
@@ 4.4 Bivariate Distributions
where $X$ and $Y$ denote the two continuous variables, whereas $\rho_{XY}$ is the correlation coefficient for $X$ and $Y$. If the real and synthetic PK target property datasets are highly similar (i.e., the synthetic dataset closely resembles the real dataset), then the absolute difference would be close to 0 or very small. Heatmap (b) in Figure 4 shows DPC on the Pearson correlation coefficient (PCC). The average DPC for PCC is 0.123. Heatmap (c) in Figure 4 shows DPC on the Spearman correlation coefficient (SCC). The average DPC for SCC is 0.138. These results indicate that the generated synthetic PK target properties resemble real data in pairwise correlations.
@@
Figure 5: MSE between real and synthetic target property data generated during training for different ablation experiments.
Abl. Exp C2 Li. Aq FS PP VD HL C.(H) C.(M) LD50 h.1 h.10 Avg
Model CBrt 0.26 0.17 0.16 0.25 0.25 0.31 0.37 0.29 0.33 0.17 0.14 0.13 0.23
DBrt 0.21 0.16 0.18 0.20 0.22 0.27 0.36 0.24 0.28 0.17 0.15 0.15 0.21
@@ 4.5 Ablation Studies
We conduct ablation studies to investigate the performance of our S2PK model given different SMILES encoders, encoder training sets, and sampling approaches for the infilling and noise model. Ablation study results (Table 4) are averages over 30 generated synthetic target property datasets, covering 90K target property values for ligands, for each ablation training run. Figure 5 graphs MSE between real and synthetic data generated during training for ablation experiments. From our ablation studies, we motivate our selected model configuration.
@@ 4.5.2 Discrete Local Gaussian Noise Model
We select different infilling strategies and noise models for ablation. Comparing noise model ablations, we measure the average MSE that each method injects into the data with Gaussian (1.19), uniform (0.53), and DLGN (0.29) ordered from most to least. This confirms that DLGN injects noise closely resembling the prior distribution. Similarly, we confirm DLGN has the best HD compared to Gaussian and uniform noise models. Comparing infilling ablations, DLGN has the best overall performance in HD and synthetic and real data MSE. This motivates the choice of DLGN for both infilling and noise models for our selected model configuration. In future work, we will explore additional noise techniques, such as quantile transformations.
@@ 5 Conclusions
The SMILES-to-Pharmacokinetic model Imagand generates synthetic PK target property data that closely resembles real data in univariate and bivariate distributions and for downstream tasks. Imagand provides a solution for the challenge of sparse overlapping PK target property data, allowing researchers to generate data to tackle complex research questions and for high-throughput screening. Future work will expand Imagand to categorical PK properties, and scale to more datasets and larger model sizes.

### v2

@@ 4.2 Data Processing
We first merge all 11 PK datasets to create a unified dataset containing 30K drugs over 12 unique PK columns for training and testing (90%/10% split) our models. Excluding the hERG dataset from which we sample 7.9K drugs, we merge the remaining 9 PK datasets for 22.1K unique drugs. We arrive at a total of 30K drugs in our unified dataset after merging the 7.9K drugs sampled from hERG into our 22.1K unique drugs from the other 9 PK datasets. We only sample 7.9K drugs from hERG to maintain balance in the unified dataset given the size imbalance of hERG compared to the other 9 PK datasets. After removing outliers ($Q1-1.5$IQR lower and $Q3+1.5$IQR upper bound), we are left with 28,397 drugs from the original 30K drugs. The 28,397 drug values for each of the 12 PK columns are then min-max scaled between the range of $[-1,1]$. Outliers are removed to ensure that Min-Max normalization does not cause unwarranted skewness in our trainset distribution, causing issues for model training. Before infilling null values using one of the average, uniform, or Gaussian distributions, or the proposed DLGN method, we store the null masks for each drug for the masked loss function.
4.3 Machine Learning Efficiency
Machine Learning Efficiency (MLE) is a measure that assesses the ability of the synthetic data to replicate a specific use case (Dankar & Ibrahim, 2021; Basri et al., 2023; Borisov et al., 2022). MLE represents the ability of the synthetic data to replace or augment real data in downstream use cases. To measure MLE, two models are trained separately using synthetic versus real data, and then their performance, measured by Mean-Squared Error (MSE), R-Squared (R2), and Pearson Correlation Coefficient (PCC), is evaluated on real data test sets and compared.
Table 2 shows the results of the PK regression tasks using real and synthetic augmented datasets. Results of these experiments suggest that a synthetic augmented dataset can outperform real data with statistical significance over many PK datasets. Additional tasks will be explored in future work. We see that synthetic data from both cGAN and Imagand can improve MLE over using only the real data. Imagand has similar or superior MLE performance compared to cGAN.
Figure 2: Synthetic PK Data Hellinger Distances (HDs).
4.4 Univariate Distributions
@@
Table 3 compares the mean and standard deviation of the real and synthetic target property values. The mean and standard deviation of the generated synthetic data closely resemble that of the real data for each PK target property. We found that normalization combined with static thresholding substantially limits the generation of invalid and out-of-range PK values. Given the underpinnings of diffusion using Gaussian reparameterization (Ho et al., 2020), diffusion methods have challenges learning and generating non-Gaussian data. This failure mode is evident in Figure 4 for non-Gaussian distributions PPBR, VDss, Half Life, and Clearance (Hep and Mic) where we see that the synthetic data fails to replicate the Log-logistic real-data distribution common in drug discovery datasets.
@@ 4.5 Bivariate Distributions
where $X$ and $Y$ denote the two continuous variables, whereas $\rho_{XY}$ is the correlation coefficient for $X$ and $Y$. If the real and synthetic PK target property datasets are highly similar (i.e., the synthetic dataset closely resembles the real dataset), then the absolute difference would be close to 0 or very small. Heatmap (b) in Figure 4 shows DPC on the Pearson correlation coefficient (PCC). The average DPC for PCC is 0.123. Heatmap (c) in Figure 4 shows DPC on the Spearman correlation coefficient (SCC). cGAN and Syngand produce synthetic data with worse DPCs compared to Imagand. The average DPC for PCC for cGAN is 0.170, and 0.154 for Syngand, compared to 0.123 for Imagand. The average DPC for SCC for cGAN is 0.187, and 0.161 for Syngand, compared to 0.138 for Imagand. These results indicate that the generated synthetic PK target properties resemble real data in pairwise correlations.
@@
Figure 5: Mean Squared Error between real and synthetic target property data generated during training for different ablation experiments.
Ablation Exp C2 Li. Aq FS PP VD HL C.(H) C.(M) LD50 h.1 h.10 Avg
Pre. CBert 0.26 0.17 0.16 0.25 0.25 0.31 0.37 0.29 0.33 0.17 0.14 0.13 0.23
DBert 0.21 0.16 0.18 0.20 0.22 0.27 0.36 0.24 0.28 0.17 0.15 0.15 0.21
@@ 4.6 Ablation Studies
We conduct ablation studies to investigate the performance of our S2PK model given different SMILES encoders, encoder training sets, and sampling approaches for the infilling and noise model. Ablation study results (Table 4) are averages over 30 generated synthetic target property datasets, covering 90K target property values for ligands, for each ablation training run. From our ablation studies we find that Imagand generates more realistic synthetic data compared to cGAN and Syngand baselines in terms of univariate distributions. Figure 5 graphs MSE between real and synthetic data generated during training for ablation experiments. From our ablation studies, we motivate our selected model configuration.
@@ 4.6.2 Discrete Local Gaussian Noise Model
We select different infilling strategies and noise models for ablation. Comparing noise model ablations, we measure the average MSE that each method injects into the data with Gaussian (1.19), uniform (0.53), and DLGN (0.29) ordered from most to least. This confirms that DLGN injects noise closely resembling the prior distribution. Similarly, we confirm DLGN has the best HD compared to Gaussian and uniform noise models. Comparing infilling ablations, DLGN has the best overall performance in HD and synthetic and real data MSE. This motivates the choice of DLGN for both infilling and noise models for our selected model configuration.
5 Discussions
Our work is a major step towards building a new class of foundational models for drug discovery trained over a diverse range of datasets. Given the problem of data sparsity, Imagand can be utilized primarily as a in silico pre-clinical tool, aimed to reduce the costs of in vitro experiments and high-throughput screening. As a research tool, scientists can utilize our models to investigate and generate properties for novel molecules to be used for downstream PBPK simulations without costly assays. Even as an initial step, Imagand has many real-world pre-clinical applications where data sparsity and data scarcity are challenges.
Although we cover a wide variety of ADMET datasets, most of these datasets are in vitro. One of the critical challenges in drug discovery is quantitative in vitro-to-in vivo extrapolation (QIVIVE). QIVIVE is an approach that extrapolates from in vitro concentration-response data to in vivo safe exposures or to identify exposure levels causing adverse effects. For future work, we will look to extend our model to include in vivo datasets and to investigate new applications of Imagand for QIVIVE. Future work will look to further explore DLGN as an alternative to Gaussian noise, including implementing the derived formulation, conducting scaling law analysis, and to compare against a broad range of models and benchmarks.
@@ 6 Conclusions
The SMILES-to-Pharmacokinetic model Imagand generates synthetic PK target property data that closely resembles real data in univariate and bivariate distributions and for downstream tasks. Imagand provides a solution for the challenge of sparse overlapping PK target property data, allowing researchers to generate data to tackle complex research questions and for high-throughput screening. Future work will expand Imagand to categorical PK properties, and scale to more datasets and larger model sizes. In future work we will look to explore additional reparameterization tricks for diffusion, such as discrete diffusion (Austin et al., 2021), to extend our methodology to be capable of learning and generating synthetic data following categorical and Log-logistic distributions common in drug discovery datasets.
Appendix A Discrete Local Gaussian Sampling
Theorem A.1.
Let $D_{KL}(P||Q)$ be the Kullback-Leibler (KL) divergence for a model probability distribution $Q$ and the true probability distribution $P$. We then have:
$lim_{N\to\infty}D_{KL}(X||\mathbb{\phi}(X=x))=lim_{N\to\infty}\sum_{\chi}Xlog(\frac{X}{\mathbb{\phi}(X=x)})=0$
Proof.
Since $F_{X}^{-1}=lim_{N\to\infty}\hat{F}_{X_{N}}^{-1}$, and as the standard deviation $\frac{\sigma}{N}\to 0$ as $N\to\infty$, then $\mathbb{\phi}(X)\sim F_{X}^{-1}$, as $N\to\infty$. Since $F_{X}^{-1}=X$, then $D_{KL}(X||\mathbb{\phi}(X=x))=0$ as $N\to\infty$. ∎
From Theorem A.1 we see that DLG sampling becomes increasingly similar to the real data distribution as the number of bins increases. Empirically, using DLG sampling results in smoother training dynamics compared to Gaussian noise as well as higher-quality generated data. Theoretically, this may be because each de-noising step is smaller, which we see empirically, in turn making them easier to learn, when the noise is increasingly similar to the original distribution, especially given skewed or multi-modal real distributions. DLG noise also enables new training modalities, which we plan on exploring in future work; bypassing latent model training, and when prior real data distributions are known and well-defined.
Appendix B Training Details
Imagand Model Diffusion Training
Layers 12 Learning Rate 1e-3
Heads 16 Weight Decay 5e-2
MLP Dim. 768 Epoch 3000
Emb. Dropout 10% Batch Size 256
Num Patches 48 Warmup 200
Cond. Emb. Size 768 Timesteps (Train) 2000
Time Emb. Size 64 Timesteps (Infer.) 150
PK Emb. Size 256 EMA Gamma ($\gamma$) 0.994
B.1 Static Thresholding
We apply elementwise clipping the PK predictions to $[-1,1]$ as static thresholding, similar to Saharia et al. (2022); Ho et al. (2020). Since PK data is min-max scaled to the same $[-1,1]$ range as a preprocessing step, static thresholding is essential to prevent the generation of invalid and out-of-range PK values.
Appendix C Convergence Analysis
Recent work by (Nakano, 2025) has proved the convergence of the original version of DDPM models as defined in (Ho et al., 2020). As our work is based on the original version of DDPM with only a modification to the noise used, we leave the majority of the proof of convergence to the work by (Nakano, 2025). Instead, to properly apply (Nakano, 2025) proof of convergence of DDPM to our formulation, we show how our modifications of DLG noise respect key conditions and lemmas required by (Nakano, 2025). Relating to our noise modification is condition (H3) in (Nakano, 2025) proof of convergence
Condition C.1 (H3 in (Nakano, 2025)).
Let $z_{i}$ be the denoising term in a diffusion process. The function $z_{i}$ for the noise estimation satisfies
$\underset{i=1,...,n}{max}||z_{i}||_{\infty}=O((log\ n)^{\kappa_{1}}),\quad n\to\infty$
for some constant $\kappa_{1}>0$
Proof.
Utilizing conditions (H1), (H2), and Lemma 2 from (Nakano, 2025), we have
$||\triangledown logp_{i}||_{\infty}\leq C_{0}/\sqrt{\bar{\alpha}_{i}}\leq C_{0}(log\ n)^{\kappa_{1}}$
Given the simplified version of the diffusion objective (as derived by (Ho et al., 2020)), it is known that the objective is equivalent to the score-matching objective (Chen et al., 2023). More precisely
$s_{i}(x):=-\frac{1}{\sqrt{1-\bar{\alpha}_{i}}}z_{i}(x)$
and the score function $\triangledown logp_{i}(\cdot)$ of $x_{i}$, $i=1,...,n$,
$\mathbb{E}|s_{i}(x_{i})-\triangledown logp_{i}(x_{i})|^{2}$
Hence, it is natural to assume that the norm of the estimated score function $s_{i}$ is bounded by $C^{\prime}_{0}(log\ n)^{\kappa}$ with some $C^{\prime}_{0}$. Given the following definition for DLG $z_{i}$
$z_{i}=\mathcal{N}(x;\hat{F}_{X_{N}}^{-1}(U),\frac{\phi}{N^{2}}\mathbf{I})$
where $z_{i}\sim F_{X}^{-1}=p_{data}$ leads to the condition (H3). ∎
Appendix D Comparison to Baseline
We compare Imagand with a baseline in Conditional GAN (cGAN) (Mirza & Osindero, 2014) with 1.8M parameters and Syngand (Hu et al., 2024) with 9M parameters. Similar to in Imagand, SMILES-embeddings from a pre-trained T5 model are used conditionally by the cGAN model to generate PK properties as output for a specific drug. Compared to earlier results, Table 6, Figure 7, and Figure 6 shows that Imagand is able to generate more realistic synthetic data compared to cGAN and Syngand.
Mean Std
Data Real Imgd cGAN Sygd Real Imgd cGAN Sygd
Caco2 0.118 0.137 0.144 0.582 0.388 0.375 0.271 0.120
Lipophilicity 0.184 0.179 0.200 0.615 0.417 0.386 0.298 0.185
AqSolDB 0.106 0.107 0.132 0.102 0.412 0.362 0.291 0.182
FreeSolv 0.103 0.123 0.099 0.267 0.421 0.400 0.294 0.134
PPBR 0.570 0.562 0.628 0.974 0.496 0.467 0.367 0.091
VDss -0.603 -0.615 -0.661 -0.985 0.442 0.389 0.310 0.065
Half life -0.557 -0.559 -0.614 -0.982 0.450 0.419 0.315 0.058
Clearance (H) -0.549 -0.559 -0.593 -0.974 0.605 0.551 0.465 0.086
Clearance (M) -0.670 -0.676 -0.738 -0.985 0.445 0.382 0.312 0.064
LD50 -0.038 -0.054 -0.044 0.043 0.372 0.331 0.272 0.153
hERG 1uM 0.036 0.031 0.048 0.115 0.338 0.319 0.252 0.325
hERG 10uM 0.027 0.030 0.018 0.073 0.335 0.316 0.246 0.304
Figure 6: Distributions of ligand PK properties and synthetic PK Data Hellinger Distances (HDs) for cGAN. Blue, synthetic distributions; orange, real distributions.
Figure 7: Distributions of ligand PK properties (log-scale) and synthetic PK Data Hellinger Distances (HDs) for Syngand. Blue, synthetic distributions; orange, real distributions.